A machine can be purchased for $60,000 and used for five years, yielding the following net incomes. In projecting net incomes, straight-line depreciation is applied, using a five-year life and a zero salvage value.

Year 1 Year 2 Year 3 Year 4 Year 5
Net income $ 3,900 $ 9,900 $ 32,000 $ 14,700 $ 39,600


Compute the machine’s payback period (ignore taxes). (Round your intermediate calculations to 3 decimal places and round payback period answer to 3 decimal places.)

Year Net Income Depreciation Net Cash Flow Cumulative Cash Flow
0 (60,000) $(60,000)
1 $3,900
2 9,900
3 32,000
4 14,700
5 39,600
Payback period =
Answer #1

Depreciation = 60000 / 5 = 12000

Year Net Income Depreciation Net Cash Flow Cumulative Cash Flow
0 (60,000) $(60,000)             
1 $3,900 12000 15900 (44100)             (Add 15900 to (60000))
2 9,900 12000 21900 (22200)              (Add 21900 to (44100))
3 32,000 12000 44000 21800               (Add 44000 to (22200))
4 14,700 12000 26700 48500                (Add 26700 to 21800)
5 39,600 12000 51600 100100                (Add 51600 to 48500)

Payback period= A + B / C

A= last period of negative cumulative cash flow

B=Absolute value of cumulative cash flow at end of period A

C=Total cash inflow during period following A

= 2 + 22200 / 26700

   2   + 0.83

   = 2.83 Years